OpenStudy (anonymous):

Which of the following simultaneous equations is not parallel to the other 3? A. -2x = 3+y B. y = 5-2x C. x = 1/2y + 6 D. X = -1/2y

OpenStudy (anonymous):

Convert all the answer choices to slope-intercept form for direct comparison: y = mx + b. To do this you need to isolate y; for example, B is already in slope-intercept form. Your answer would be the one equation with a differing slope, since parallel lines all have the same slope.

OpenStudy (anonymous):

actually, for B to be in y=mx + b form you just need to invert the terms, so y = 5 - 2x is y = -2x +5

OpenStudy (anonymous):

Do you get how to do it?

OpenStudy (anonymous):

Yeah, the answer is b right?

OpenStudy (anonymous):

I just didn't put them onto a graph haha

OpenStudy (anonymous):

No, it isn't actually, since A is y = -2x + 3 . A and B have the same slope of -2 so neither can be your answer. All you need to do is find out whether C or D have a different slope and the answer choice that does not have a slope of -2 in y = mx + b form is your answer.

OpenStudy (anonymous):

Ok, thank you!

OpenStudy (anonymous):

No problem